# Language Models are Injective and Hence Invertible

### v3

@@ 1 Introduction
A core question in understanding large language models is whether their internal representations faithfully preserve the information in their inputs. Since Transformer architectures rely heavily on non-linearities, normalization, and many-to-one attentions mechanisms, it is often assumed that they discard information: different inputs could collapse to the same hidden state, making exact recovery of the input impossible. This view motivates concerns around transparency, robustness, and safe deployment, as it suggests that the link between text and representation is inherently lossy.
@@ 2 Transformers are injective
These facts allow us to state and prove injectivity results without relying on asymptotics.
We begin by establishing the analytic structure of the architecture.
@@
Figure 3: Seeking collisions in a large-scale prompt set (§4.1). The minimum distances between last-token states are far above the collision threshold $10^{-6}$: (left) across layers for GPT-2 and Gemma-3 families (one dot per layer), (right) across depth for GPT-2 Small, where distances grow with depth.
@@ 3 Exact prompt recovery via SipIt
In the previous section, we have proven that decoder-only Transformers are almost surely injective, i.e., different prompts map to different hidden states. We now show how this property can be used in practice to reconstruct the exact input prompt given hidden states at some layer. We call this algorithm SipIt (Sequential Inverse Prompt via ITerative updates).
Formally, recall from §2 that the mapping from a prompt $\mathrm{s}$ to its last-token state is almost surely injective. Since the last state is itself a deterministic function of the hidden matrix at any layer $\ell$, injectivity extends to the full representation
@@
Algorithm 1 Sip-It: Sequential Inverse Prompt via Iterative Updates
@@ 4.1 Searching for collisions
Figure 4: Exhaustive collision search on the 1010 closest prefix prompts. The boxplots look flat and uneventful, and that is the point: even under stress-test conditions with billions of candidate pairs, all minima stay well above the collision threshold, showing that nothing collapses.
Model L2 Distance (min)
@@
Table 1: Minimum pairwise distance between last-token states in the first, middle, and final layers of four models. All values are well above the collision threshold $10^{-6}$ (no collisions).
We observed no collisions across all models and layers: distinct prompts always yielded distinct last-token states. Figure 3 (left) shows the per-layer minimum distances for the Gemma3 pretrained (Team et al., 2025) and GPT-2 (Radford et al., 2019) families, with strictly positive values throughout. Table 1 complements this by reporting the same statistic for Llama-3.1-8B (Grattafiori et al., 2024), Mistral-7B-v0.1 (Jiang et al., 2023), Phi-4-mini-instruct (Microsoft et al., 2025) and TinyStories-33M (Eldan & Li, 2023), again showing clear separation at the first, middle, and last layers.
Finally, Figure 3 (right) zooms in on GPT-2 Small, revealing that these distances typically increase with depth. Additional results for GPT-2 Medium, GPT-2 Large and Gemma3 (1B, 4B, 12B) appear in Appendix E, confirming the same trend.
@@
This exhaustive experiment helps rule out the possibility that earlier observations were simply due to chance in random sampling rather than a true absence of collisions. While a complete search over all possible prompts would be ideal, it is computationally infeasible. The number of unique prompts grows exponentially with sequence length, and the number of pairwise comparisons grows even faster. For context, even with single-token prompts and the vocabulary size of Gemma3-1B, there are already over 34 trillion possible prompt pairs, making exhaustive evaluation entirely impractical. Our compromise still revealed structure: we identified 5 prompt pairs with highly similar last-token embeddings, suggesting overlapping semantic content and motivating us to ask whether distinct next tokens could preserve meaning, i.e., yield essentially identical last-token hidden states.
Figure 4 reports the resulting distributions (min/median/mean/max) as boxplots for both GPT-2 Small and Gemma3-1B, with distances far from zero (no collision), confirming local injectivity as predicted by our theory.
@@ 4.2 Invertibility results
Table 2: Prompt inversion: SipIt ensures exact recovery efficiently, unlike HardPrompts (no recovery) or brute force (infeasible runtimes).
We now test whether the theoretical injectivity translates into exact recovery on pre-trained models. Using SipIt with only the hidden states at a fixed layer, we attempt to reconstruct the full prompt token-by-token for GPT-2 Small. We sample 100 prompts, with a $90\%$-$10\%$ split between meaningful sentences and random token sequences (to test robustness in unstructured cases), and attempt to reconstruct them from hidden states.
We compare against HardPrompts (Wen et al., 2023), which leverages gradient signals for approximate prompt discovery, and against a SipIt ablation without the gradient-guided candidate policy (BruteForce).
[…]
Results are reported in Table 2. Across all prompts (20 tokens each), SipIt recovers the exact sequence with $100\%$ token-level accuracy (no errors, no collisions), matching the theoretical guarantee of linear-time convergence.
In contrast, HardPrompts fails to recover the true input in most cases, while BruteForce eventually succeeds but at a prohibitive computational cost, requiring several orders of magnitude longer.
@@ 5 Related work
Viewed as functions on $\mathbb{R}^{d}$, individual Transformer components are clearly non-injective: LayerNorm collapses along per-example statistics (Ba et al., 2016), residual connections can cancel, and in attention-only stacks, rank decays doubly-exponentially with depth (Dong et al., 2021). Likewise, on the output side, the softmax bottleneck constrains the distributions reachable by language models (Yang et al., 2018). From this algebraic perspective, Transformers seem inherently many-to-one, while in a generative sense, they can also behave one-to-many when different prompts lead to the same continuation.
Our focus is different: we study the discrete-to-continuous map from prompts $\mathrm{s}\in\mathcal{V}^{\leq K}$ to hidden states in $\mathbb{R}^{d}$. In this setting, analytic viewpoints on Transformer computation become powerful: treating each layer as a real-analytic map yields almost-sure guarantees that hold at finite width, depth, and training horizon. Recent work has adopted this angle for related properties: Jiang & Haghtalab (2025) show that building blocks of modern architectures are almost always surjective, while Sutter et al. (2025) prove that Transformers at random initialization are almost surely injective with respect to the entire hidden-state matrix (and only at initialization).
Differently, we prove injectivity with respect to the parameters and at the task-relevant last-token state; crucially, we show that injectivity is not an initialization artifact but persists under training.
@@
Inverse problems seek to recover an unknown input $x$ from observations $y$ produced by a forward process $y=f(x)$ (Sun et al., 2021). Within this landscape, language model inversion asks whether one can reconstruct a model’s input prompt from outputs or internal signals.
Several approaches have explored this idea. Output-to-prompt methods infer prompts from generated continuations, yielding approximate reconstructions that are often semantically similar rather than exact (Zhang et al., 2024). Recent work by Morris and coauthors shows that model outputs are information-rich even in black-box settings: Morris et al. (2023b) train a separate inverter to map next-token probability vectors to text, and Nazir et al. (2025) extend this by taking sequences of logprobs, applying a linear compression to embedding dimension, and training an encoder-decoder inverter; this achieves higher exact-match rates but still without guarantees. Complementarily, Morris et al. (2023a) reconstruct text from encoder logits via a trained iterative inverter. These contributions highlight privacy risks when probabilities or embeddings are exposed, but they differ from our setting: they rely on trained inverters, remain approximate, and do not invert hidden states of decoder-only LMs.
A related line of work frames the task as automated prompt optimization, casting prompt design as discrete sequence optimization aligned with downstream performance (Guo et al., 2025; Sun et al., 2022; Deng et al., 2022); methods such as AutoPrompt (Shin et al., 2020) and Hard Prompts Made Easy (Wen et al., 2023) use gradient signals to discover effective, but approximate, prompts.
Unlike prior work, which yields approximate reconstructions from outputs, logits, or logprobs, our approach is training-free, efficient, and comes with provable linear-time guarantees for exact recovery from internal states.
@@ 6 Discussion and conclusions
Beyond theory, the findings carry practical and legal implications. Hidden states are not abstractions but the prompt in disguise. Any system that stores or transmits them is effectively handling user text itself. This affects privacy, deletion, and compliance: even after prompt deletion, embeddings retain the content. Regulators have sometimes argued otherwise; for example, the Hamburg Data Protection Commissioner claimed that weights do not qualify as personal data since training examples cannot be trivially reconstructed (HmbBfDI, 2024). Our results show that at inference time user inputs remain fully recoverable. There is no “free privacy” once data enters a Transformer.
@@
We provide complete resources to ensure reproducibility of our results. The assumptions, definitions, and full proofs can be found in section 2 and appendices A, C and B (analytic tools and model specification in appendices A and B; almost-sure injectivity and preservation under training in appendix C; SIP-IT correctness, verifier, and margin analysis in appendix D). Implementation details for SIP-IT, including pseudocode, are provided in section 3 and algorithm 1 and further elaborated in appendix E. Our experimental setup (hardware and software versions) is described in section 4, while dataset details and the prompt-sampling procedure for the 100k-prompt benchmark are given in section 4.1. Finally, the supplementary materials include an anonymized code repository with end-to-end scripts, fixed seeds, configuration files, and a comprehensive README with step-by-step reproduction instructions.
@@ Appendix A Preliminaries
This section fixes notation the notation used throughout the main paper and the appendix (subsection A.1), and it introduces real-analyticity as the organizing theme (subsection A.2). We first review the vector-space notion and its basic closure/composition properties (subsubsection A.2.1), together with a zero-set principle used in measure-zero arguments. We then extend these ideas to maps between matrix spaces (subsubsection A.2.2) via vectorization/matricization and note that analyticity is preserved under matrix compositions. To streamline later proofs, we summarize real-analytic building blocks commonly used in transformer layers–polynomials, exponential/logarithm, softmax, row normalization, matrix products, Hadamard scaling, and stacking (subsubsection A.2.3). Finally, in subsection A.3, we collect differential and topological tools–Fréchet derivatives and the Hessian, standard facts on $\mathbb{R}^{p}$, the inverse function theorem, and pushforwards/absolute continuity–which we use for local invertibility and absolute-continuity arguments. Readers already comfortable with these topics can skim now and return to specific subsections as needed.

### v4

@@ 1 Introduction
A core question in understanding large language models is whether their internal representations faithfully preserve the information in their inputs. Since Transformer architectures rely heavily on non-linearities, normalization, and many-to-one attention mechanisms, it is often assumed that they discard information: different inputs could collapse to the same hidden state, making exact recovery of the input impossible. This view motivates concerns around transparency, robustness, and safe deployment, as it suggests that the link between text and representation is inherently lossy.
@@ 2 Transformers are injective
These facts allow us to state and prove injectivity results without relying on asymptotics. We begin by establishing the analytic structure of the architecture.
@@
Figure 3: Seeking collisions in a large-scale prompt set: The minimum distances between last-token states are far above the collision threshold $10^{-6}$: (left) across layers for GPT-2 and Gemma-3 families (one dot per layer), (right) across depth for GPT-2 Small; distances grow with depth.
@@ 3 Exact prompt recovery via SipIt
In the previous section, we have proven that decoder-only Transformers are almost surely injective, i.e., different prompts map to different hidden states. We now show how this property can be used in practice to reconstruct the exact input prompt given hidden states at some layer. We call this algorithm SipIt (Sequential Inverse Prompt via ITerative updates). [^3] Implementation available at https://github.com/giorgosnikolaou/SIPIT.
Threat model.
This paper focuses on the injectivity result and its algorithmic consequence; we do not define a full adversarial model. A natural setting where SipIt applies is one in which an adversary obtains the hidden-state sequence—for instance, through a leaked KV-cache, a shared-inference pipeline, or an API that exposes intermediate representations. Our injectivity result guarantees that exact recovery from only the final embedding is possible in principle, but designing an efficient algorithm for that setting is nontrivial and left to future work; here we assume access to all per-position states at a given layer $\ell$.
Recall from §2 that the mapping from a prompt $\mathrm{s}$ to its last-token state is almost surely injective. Since the last state is itself a deterministic function of the hidden matrix at any layer $\ell$, injectivity extends to the full representation
@@
Algorithm 1 SipIt: Sequential Inverse Prompt via ITerative updates
@@
Theorem 3.2 (Robustness of SipIt).
Under the assumptions of Theorem 2.3, fix a layer $\ell$ and define, for any prefix $\pi$ and time $t$,
$\Delta_{\pi,t}\ :=\ \min_{v\neq v^{\prime}\in\mathcal{V}}\big\|\mathbf{h}_{t}(\pi\oplus v)-\mathbf{h}_{t}(\pi\oplus v^{\prime})\big\|_{2}.$
Let $\mathrm{s}=\langle\mathrm{s}_{1},\ldots,\mathrm{s}_{T}\rangle$, define the prefixes $\pi_{t}=\langle\mathrm{s}_{1},\ldots,\mathrm{s}_{t-1}\rangle$ and suppose access to the perturbed hidden states
$\widehat{\mathbf{h}}_{t}(\pi_{t}\oplus\mathrm{s}_{t})=\mathbf{h}_{t}(\pi_{t}\oplus\mathrm{s}_{t})+\mathbf{e}_{t},\qquad\|\mathbf{e}_{t}\|_{2}<\tfrac{\Delta_{\pi_{t},t}}{2},\quad t\in[T].$
Then SipIt recovers the true sequence $\mathrm{s}$ with probability one, and terminates in at most $T|\mathcal{V}|$ steps.
Proof in Appendix D, Thm. D.2, Prop. D.2.
@@ 4.1 Searching for collisions
Figure 4: Exhaustive collision search on the $10$ closest prefix prompts. The boxplots look flat and uneventful, and that is the point: even under stress-test conditions with billions of candidate pairs, all minima stay well above the collision threshold, showing that nothing collapses.
Model $\bm{\ell}_{\mathbf{2}}$ Distance (min)
@@
Table 1: Minimum pairwise distance between last-token states in the first, middle, and final layers of four models. All values are well above the collision threshold $10^{-6}$.
We observed no collisions across all models and layers: distinct prompts always yielded distinct last-token states. Figure 3 (left) shows the per-layer minimum distances for the Gemma3 pretrained (Team et al., 2025) and GPT-2 (Radford et al., 2019) families, with strictly positive values throughout. Table 1 complements this by reporting the same statistic for Llama-3.1-8B (Grattafiori et al., 2024), Mistral-7B-v0.1 (Jiang et al., 2023), Phi-4-mini-instruct (Microsoft et al., 2025) and TinyStories-33M (Eldan and Li, 2023), again showing clear separation at the first, middle, and last layers. Finally, Figure 3 (right) zooms in on GPT-2 Small, revealing that these distances typically increase with depth. Additional results for GPT-2 Medium, GPT-2 Large and Gemma3 (1B, 4B, 12B) appear in Appendix E, confirming the same trend.
Model $\bm{\ell}_{\mathbf{2}}$ Distance (min)
FP4 INT8 FP32
Llama-3.1-8B 2.281 6.597 1.274
Mistral-7B-v0.1 1.748 2.692 1.136
Phi-4-mini-instruct 18.368 20.956 8.780
Table 2: Quantized Models: Minimum pairwise distance between last-token states in the final layer of three quantized models.
Model Size $\bm{\ell}_{\mathbf{2}}$ Distance (min)
layer 1 layer $L/2$ layer $L$
phi-4 14B 0.010 1.025 8.759
Llama-3.1-70B 70B 0.005 0.465 3.975
Table 3: Large Models: Minimum pairwise distance between last-token states in the first, middle, and final layers of two large models.
@@
[…]
This exhaustive experiment helps rule out the possibility that earlier observations were simply due to chance in random sampling rather than a true absence of collisions. While a complete search over all possible prompts would be ideal, it is computationally infeasible. The number of unique prompts grows exponentially with sequence length, and the number of pairwise comparisons grows even faster. For context, even with single-token prompts and the vocabulary size of Gemma3-1B, there are already over 34 billion possible prompt pairs, making exhaustive evaluation entirely impractical. Our compromise still revealed structure: we identified 5 prompt pairs with highly similar last-token embeddings, suggesting overlapping semantic content and motivating us to ask whether distinct next tokens could preserve meaning, i.e., yield essentially identical last-token hidden states.
Figure 4 reports the resulting distributions as boxplots for both GPT-2 Small and Gemma3-1B, with distances far from zero (no collision), confirming local injectivity as predicted by our theory.
FP4 and INT8 weight quantization.
To assess how weight quantization affects pairwise representation distances, we conducted additional experiments with FP4 and INT8 quantization on several models (Llama-3.1-8B, Phi-4-mini-instruct, and Mistral-7B-v0.1). We further extended this analysis to FP4-quantized 14B and 70B parameter models, namely Phi-4 (14B) and Llama-3.1-70B. As shown in Tables 3 and 3, across all tested models quantization (1) does not introduce any collisions, (2) more than doubles the minimum distance between representations, thereby preserving the integrity of the representation space, and (3) maintains this separation even as model size increases substantially.
@@ 4.2 Invertibility results
We now test whether the theoretical injectivity translates into exact recovery on pre-trained models. Using SipIt with only the hidden states at a fixed layer, we attempt to reconstruct the full prompt token-by-token for GPT-2 Small. We sample 100 prompts, with a $90\%$-$10\%$ split between meaningful sentences and random token sequences (to test robustness in unstructured cases), and attempt to reconstruct them from hidden states. We compare against HardPrompts (Wen et al., 2023), which leverages gradient signals for approximate prompt discovery, and against a SipIt ablation that replaces the gradient-guided candidate policy with the uniformly random policy (BruteForce).
Model Vocab size Inversion Performance
Accuracy Time (s) Vocab explored (%)
Mistral-7B-v0.1 32000 100% 111.78 $\pm\,\,\,$ 46.50 0.19 $\pm$ 0.08 %
Llama-3.1-8B 128255 100% 549.48 $\pm$ 265.75 0.21 $\pm$ 0.10 %
Table 4: Inversion performance on FP4-quantized models with different vocabulary sizes. SipIt recovers all tokens with 100% accuracy while exploring less than $0.22\%$ of the vocabulary on average.
@@
Table 5: SipIt ensures efficient exact recovery, unlike HardPrompts (no recovery) or BruteForce (infeasible runtime).
Results are reported in Table 5. Across all prompts (20 tokens each), SipIt recovers the exact sequence with $100\%$ token-level accuracy (no errors, no collisions), matching the theoretical guarantee of linear-time convergence. In contrast, HardPrompts completely fails to recover the input, while BruteForce eventually succeeds but at a prohibitive computational cost, requiring several orders of magnitude longer.
Robustness and vocabulary scaling.
The theoretical analysis in Theorem 3.2 shows that our inversion algorithm is robust to a certain level of noise while maintaining linear scaling in vocabulary size. To empirically validate this, we use FP4-quantized versions of Mistral-7B-v0.1 ($\approx 32\text{K}$ vocabulary size) and Llama-3.1-8B ($\approx 128\text{K}$). We sample 50 prompts (10 tokens each) and attempt to reconstruct them from hidden states corrupted by FP4 weight quantization. As shown in Table 4, SipIt reconstructs all inputs with perfect accuracy while exploring, on average, less than $0.22\%$ of the vocabulary, demonstrating that the gradient-based heuristic is both robust to quantization noise and highly efficient. From a complexity perspective, the nearly constant percentage of tokens explored across the two vocabulary scales empirically confirms the predicted linear scaling.
Effect of layer depth.
@@ 5 Related work
Viewed as functions on $\mathbb{R}^{d}$, individual Transformer components are clearly non-injective: LayerNorm collapses along per-example statistics (Ba et al., 2016), residual connections can cancel, and in attention-only stacks, rank decays doubly-exponentially with depth (Dong et al., 2021). Likewise, on the output side, the softmax bottleneck constrains the distributions reachable by language models (Yang et al., 2018). From this algebraic perspective, Transformers seem inherently many-to-one, an intuition echoed by classical completeness and universal-approximation theorems for Transformers, which show that highly many-to-one maps can be represented in principle; we briefly review these results in Appendix F.
Our focus is different: we study the discrete-to-continuous map from prompts $\mathrm{s}\in\mathcal{V}^{\leq K}$ to hidden states in $\mathbb{R}^{d}$. In this setting, analytic viewpoints on Transformer computation become powerful: treating each layer as a real-analytic map yields almost-sure guarantees that hold at finite width, depth, and training horizon (Appendix F surveys which modern LLMs satisfy this assumption and proves the analyticity for all activation functions encountered in practice). Recent work has adopted this angle for related properties: Jiang and Haghtalab (2025) show that building blocks of modern architectures are almost always surjective, while Sutter et al. (2025) prove that Transformers at random initialization are almost surely injective with respect to the entire hidden-state matrix (and only at initialization). Differently, we prove injectivity with respect to the parameters and at the task-relevant last-token state; crucially, we show that injectivity is not an initialization artifact but persists under training.
@@
Model inversion asks whether one can reconstruct a model’s input prompt from outputs or internal signals (Sun et al., 2021). In the context of language models, this question has motivated a growing body of work exploring practical inversion strategies. Output-to-prompt methods infer prompts from generated continuations but yield only approximate reconstructions (Zhang et al., 2024). Recent work shows that even black-box outputs are information-rich: Morris et al. (2023b) train a separate inverter to map next-token probability vectors to text, and Nazir et al. (2025) extend this by taking sequences of logprobs, applying a linear compression to embedding dimension, and training an encoder-decoder inverter; this achieves higher exact-match rates but still without guarantees. Complementarily, Morris et al. (2023a) reconstruct text from encoder logits via a trained iterative inverter. These contributions highlight privacy risks when probabilities or embeddings are exposed, but they differ from our setting: they rely on trained inverters, remain approximate, and do not invert hidden states of decoder-only LMs.
A related line of work frames the task as automated prompt optimization, casting prompt design as discrete sequence optimization aligned with downstream performance (Guo et al., 2025; Sun et al., 2022; Deng et al., 2022); methods such as AutoPrompt (Shin et al., 2020) and Hard Prompts Made Easy (Wen et al., 2023) use gradient signals to discover effective, but approximate, prompts. Most closely related to ours, Thomas et al. (2025) recover prompts from hidden states via a sequential algorithm that uses an LLM-based policy to rank candidates; lacking injectivity guarantees, however, it must score all vocabulary tokens before committing, with no formal exactness guarantees.
Unlike all prior work, our approach is training-free, efficient, and comes with provable linear-time guarantees for exact recovery from internal states.
@@ 6 Discussion and conclusions
Beyond theory, the findings carry practical and legal implications. Hidden states are not abstractions but the prompt in disguise: any system that stores or transmits them is effectively handling user text itself, with direct consequences for privacy, deletion, and compliance (Miranda et al., 2025). The evolving regulatory landscape has not yet fully reckoned with this fact. The Hamburg Data Protection Commissioner, for instance, argued that LLM parameters do not constitute personal data because training data is transformed into abstract mathematical representations during learning, and that it “remains doubtful whether any extractable data records constitute personal data” (HmbBfDI, 2024). That analysis, however, concerns training data encoded in model weights; it does not address the hidden representations computed at inference time. Our results reveal that these representations are lossless encodings of the user’s exact input, recoverable in full via SipIt. Consequently, any system that stores, caches, or transmits hidden states is effectively handling the user’s verbatim text, and the corresponding pipelines should be subject to the same data-protection obligations as the raw prompts they encode.
@@
The experimental setup (hardware, software versions, and dataset construction) is described in §4; the 100k-prompt benchmark uses uniform sampling from four public datasets detailed in §4.1. On the theory side, every theorem stated in the main text is accompanied by a complete proof in the appendix: analytic preliminaries in Appendix A, the formal definition of the Transformer language model and the proof that it is real-analytic in Appendix B, almost-sure injectivity and its preservation under training in Appendix C, and SipIt correctness and robustness in Appendix D. Appendix E provides Implementation Details and Additional Experiments, and Appendix F verifies that all activation functions used in all modern LLMs satisfy the analyticity assumption.
@@ Acknowledgments
This work has been supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program, the MUR FIS2 grant n. FIS-2023-00942 “NEXUS” (cup B53C25001030001), and partly by Sapienza University of Rome via the Seed of ERC grant “MINT.AI” (cup B83C25001040001).
Appendix Overview
@@ Appendix A Preliminaries
This section fixes the notation used throughout the main paper and the appendix (subsection A.1), and it introduces real-analyticity as the organizing theme (subsection A.2). We first review the vector-space notion and its basic closure/composition properties (subsubsection A.2.1), together with a zero-set principle used in measure-zero arguments. We then extend these ideas to maps between matrix spaces (subsubsection A.2.2) via vectorization/matricization and note that analyticity is preserved under matrix compositions. To streamline later proofs, we summarize real-analytic building blocks commonly used in transformer layers–polynomials, exponential/logarithm, softmax, row normalization, matrix products, Hadamard scaling, and stacking (subsubsection A.2.3). Finally, in subsection A.3, we collect differential and topological tools–Fréchet derivatives and the Hessian, standard facts on $\mathbb{R}^{p}$, the inverse function theorem, and pushforwards/absolute continuity–which we use for local invertibility and absolute-continuity arguments. Readers already comfortable with these topics can skim now and return to specific subsections as needed.